# Implicit Safe Set Algorithm for Provably Safe Reinforcement Learning

### v1

@@ 4 Implicit Safe Set Algorithm for Systems with Negligible Sampling Time
4.1 Synthesize Safety Index for Continuous-time System
@@
These assumptions are easy to meet in practice. The first assumption ensures that the discrete-time approximation error is negligible, i.e. continuous-time system. The second assumption enables safety index design rule applicable with multiple moving obstacles.
Following the rules in (?), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$,
$\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i},$ (3)
where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$. Our goal is to choose these parameters such that $\mathcal{U}_{S}^{D}(x)$ is always nonempty. Under the above assumptions, the safety index design rule is constructed below. The main idea of designing the safety index is to derive the sufficient condition of parameters of $\phi$ to ensure there exist safe controls for all states.
@@
$\frac{n(\sigma+d_{min}^{n}+kv_{max})^{\frac{n-1}{n}}}{k}\leq\frac{-a_{min}}{v_{max}},$ (4)
where $v_{max}$ is the maximum relative velocity that the robot can achieve in the obstacle frame. We will show that the safety index design rule is valid to guarantee non-empty $\mathcal{U}_{S}^{D}(x)$ for continuous-time systems in Proposition 1.
4.2 Sample-Efficient Black-Box Constrained Optimization
@@
$\begin{split}&\min_{u_{t}\in\mathcal{U}}\|u_{t}-u^{r}_{t}\|^{2}\\ &\text{s.t. }\phi(f(x_{t},u_{t}))\leq\max\{\phi(x_{t})-\eta,0\}\end{split}$ (5)
Since the objective of (5) is convex, its optimal solution will always lie on the boundary of $\mathcal{U}_{S}^{D}(x)$ if $u^{r}\notin\mathcal{U}_{S}^{D}(x)$. Therefore, it is desired to have an efficient algorithm to find the safe controls on the boundary of $\mathcal{U}_{S}^{D}(x)$. To efficiently perform this black-box optimization, we propose a sample-efficient boundary approximation algorithm called Adaptive Momentum Boundary Approximation Algorithm (AdamBA), which is summarized in Algorithm 1. AdamBA leverages the idea of adaptive line search with exponential outreach/decay to efficiently search for the boundary points. The inputs for AdamBA are the approximation error bound ($\epsilon$), learning rate ($\beta$), reference control ($u^{r}$), gradient vector covariance ($\Sigma$), gradient vector number ($n$), reference gradient vector ($\vec{v}^{r}$), safety status of reference control ($S$), and the desired safety status of control solution ($S_{goal}$).
We illustrate the main boundary approximation procedures of AdamBA in Figure 2, where AdamBA is supposed to find the boundary points of $\mathcal{U}_{S}^{D}(x)$ (green area) with respect to the reference control $u^{r}\notin\mathcal{U}_{S}^{D}(x)$ (red star). The core idea of the AdamBA follows the adaptive line search (?), where three main procedures are included. I. AdamBA first initializes several unit gradient vectors (green vectors) to be the sampling directions, as shown in Figure 2(a). II. AdamBA enters the exponential outreach phase by exponential increasing the gradient vector length until they reach $\mathcal{U}_{S}^{D}(x)$ as shown in Figure 2(b). Note that we discard those gradient vectors that go out of control space (red vectors). III. Next, AdamBA enters the exponential decay phase by iteratively applying Bisection search to find boundary points as shown in Figure 2(c). Finally, a set of boundary points will be returned after AdamBA converges as shown in Figure 2(d). Note that AdamBA and the line search methods are fundamentally similar to each other, while the purpose of AdamBA is to find the boundary of safe/unsafe action, while the line search methods are to find the minimum of a function. It can be shown that the boundary approximation error can be upper bounded within arbitrary resolution.
(a) Initialize gradient vectors
@@
Figure 2: Illustration of the procedure of the AdamBA algorithm.
@@
4.3 Implicit Safe Set Algorithm for Continuous-Time System
@@
Leveraging AdamBA and the safety index design rule, we construct the implicit safe set algorithm (ISSA) to safeguard the unsafe reference control. The proposed ISSA is summarized in Algorithm 2. Note that ISSA is presented under the context of discrete-time system with negligible sampling time, and $\mathcal{U}_{S}^{D}$ (line 3 of Algorithm 2) is the same as $\mathcal{U}_{S}$ when the sampling time is negligible. The key idea of ISSA is to use AdamBA for efficient search and grid sampling for worst cases. The inputs for ISSA are the approximation error bound ($\epsilon$), the learning rate ($\beta$), gradient vector covariance ($\Sigma$), gradient vector number ($n$) and reference unsafe control ($u^{r}$). Note that ISSA is only needed when the reference control is not safe.
ISSA contains an offline stage and an online stage. In the offline stage, we synthesize the safety index according to the design rules. There are two major phases in the online stage for solving (5). In online-phase 1, we directly use AdamBA to find the safe controls on the boundary of $\mathcal{U}_{S}^{D}(x)$, and choose the control with minimum deviation from the reference control as the final output. In the case that no safe control is returned in online-phase 1 due to sparse sampling, online-phase 2 is activated. We uniformly sample the control space and deploy AdamBA again on these samples to find the safe control on the boundary of $\mathcal{U}_{S}^{D}(x)$. It can be shown in Section 5.2 that the ISSA algorithm is guaranteed to find a feasible solution of (5) since the nonempty set of safe control is guaranteed by the safety index design rule in (4). And we show in Section 5 that solution ensures forward invariance to the set $\mathcal{S}:=\mathcal{X}_{S}\cap\{x\mid\phi(x)\leq 0\}$.
Although the ISSA algorithm builds upon the safe set algorithm (?), the proposed safety index synthesis and AdamBA algorithm can be applied to other energy function-based methods to generate safe controls with or without an explicit analytical dynamics model.
5 Theoretical Results for ISSA When the Sampling Time is Negligible
@@
If the control system satisfies the conditions in Assumption 1 and Assumption 2, and the safety index design follows the rule described in Section 4.1, the implicit safe set algorithm in Algorithm 2 guarantees the forward invariance to the set $\mathcal{S}\subseteq\mathcal{X}_{S}$.
@@
5.1 Feasibility of the Safety Index for Continuous-Time System
@@
If 1) the dynamic system satisfies the conditions in Assumption 1 and Assumption 2; and 2) the safety index is designed according to the rule in Section 4.1, then the robot system in 2D plane has nonempty set of safe control at any state, i.e., $\mathcal{U}_{S}^{D}(x)\neq\emptyset,\forall x$.
@@ 5.1.1 Preliminary Results
If the dynamic system satisfies the conditions in Assumption 1 and Assumption 2 and there is only one obstacle in the environment, then the safety index design rule in Section 4.1 ensures that $\mathcal{U}_{S}^{D}(x)\neq\emptyset$ for $x$ such that $\phi(x)-\eta\geq 0$.
@@
$\mathcal{U}_{S}^{D}(x)=\{u\in\mathcal{U}\mid\phi(f(x,u))\leq\phi(x)-\eta\}$ (6)
According to the first condition in Assumption 2, we have $dt\rightarrow 0$. Therefore, the discrete-time approximation error approaches zero, i.e., $\phi(f(x,u))=\phi(x)+dt\cdot\dot{\phi}(x,u)+\Delta$, where $\Delta\rightarrow 0$. Then we can rewrite (6) as:
$\mathcal{U}_{S}^{D}(x)=\{u\in\mathcal{U}\mid\dot{\phi}\leq-\eta/dt\}$ (7)
@@
$\forall x\text{ s.t. }\phi(x)\geq\eta,\exists u,\text{s.t. }\ddot{d}\geq\frac{\eta/dt-nd^{n-1}\dot{d}}{k}.$ (8)
Note that in the 2D problem, $\ddot{d}=-a\cos(\alpha)+v\sin{(\alpha)}w$ and $\dot{d}=-v\cos(\alpha)$. According to Assumption 1, there is a surjection from $u$ to $(a,w)\in W:=\{(a,w)\mid a_{min}\leq a\leq a_{max},w_{min}\leq w\leq w_{max}\}$. Moreover, according to (3), $\phi$ for the 2D problem only depends on $\alpha$, $v$, and $d$. Hence condition $\forall x\text{ s.t. }\phi(x)\geq\eta$ can be translated to $\forall(\alpha,v,d)\text{ s.t. }\sigma+d_{min}^{n}-d^{n}-kv\cos(\alpha)\geq\eta$. Denote the later set as
$\Phi:=\{(\alpha,v,d)\mid\sigma+d_{min}^{n}-d^{n}-kv\cos(\alpha)\geq\eta,v\in[0,v_{max}],d\geq 0,\alpha\in[0,2\pi)\}.$ (9)
@@
$\forall(\alpha,v,d)\in\Phi,\exists(a,w)\in W,\text{ s.t. }-a\cos(\alpha)+v\sin(\alpha)w$ $\geq\frac{\eta/dt+nd^{n-1}v\cos(\alpha)}{k}.$ (10)
@@
$\forall(\alpha,v,d)\in\Phi,\exists(a,w)\in W,\text{ s.t. }-\frac{a}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k}\geq 0,$ (11)
@@
$\min_{(\alpha,v,d)\in\Phi}\max_{(a,w)\in W}(-\frac{a}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k})\geq 0.$ (12)
@@
$\min_{(\alpha,v,d)\in\Phi}\max_{(a,w)\in W}(-\frac{a}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k})$ (13)
$=$ $\min_{(\alpha,v,d)\in\Phi}(-\frac{a_{min}}{v}+[\tan(\alpha)]_{+}w_{max}+[\tan(\alpha)]_{-}w_{min}-\frac{nd^{n-1}}{k})$ (14)
$=$ $\min_{\alpha\in[0,2\pi),v\in(0,v_{max}]}(-\frac{a_{min}}{v}+[\tan(\alpha)]_{+}w_{max}+[\tan(\alpha)]_{-}w_{min}-\frac{n(\sigma+d_{min}^{n}+kv\cos(\alpha))^{\frac{n-1}{n}}}{k})$ (15)
$=$ $-\frac{a_{min}}{v_{max}}-\frac{n(\sigma+d_{min}^{n}+kv_{max})^{\frac{n-1}{n}}}{k}.$ (16)
The first equality eliminates the inner maximization where $[\tan(\alpha)]_{+}:=\max\{\tan(\alpha),0\}$ and $[\tan(\alpha)]_{-}:=\min\{\tan(\alpha),0\}$. The second equality eliminates $d$ according to the constraint in $\Phi$. The third equality is achieved when $\alpha=0$ and $v=v_{max}$. According to the safety index design rule in Section 4.1, (16) is greater than or equal to zero. Hence (12) holds, which then implies that the inequality in (10) holds.
@@
If the dynamic system satisfies the contidions in Assumption 1 and Assumption 2 and there is only one obstacle in the environment, then the safety index design rule in Section 4.1 ensures that $\mathcal{U}_{S}^{D}(x)=\mathcal{U}$ for any $x$ that $\phi(x)-\eta<0$.
@@
$\mathcal{U}_{S}^{D}(x)=\{u\in\mathcal{U}\mid\phi(f(x,u))\leq 0\}$ (17)
According to the first assumption in Assumption 2, we have $dt\rightarrow 0$. Therefore, the discrete-time approximation error approaches zero, i.e., $\phi(f(x,u))=\phi(x)+dt\cdot\dot{\phi}(x,u)+\Delta$, where $\Delta\rightarrow 0$. Then we can rewrite (17) as:
$\mathcal{U}_{S}^{D}(x)=\{u\in\mathcal{U}\mid\dot{\phi}\leq-\phi/dt\}$ (18)
Note that $\eta=0$ according to the safety index design rule, then $\phi(x)-\eta<0$ implies that $\phi(x)<0$. Hence $-\phi/dt\rightarrow\infty$ since $dt\rightarrow 0$. Then as long as $\dot{\phi}$ is bounded, $\mathcal{U}_{S}^{D}(x)=\mathcal{U}$.
Now we show that $\dot{\phi}$ is bounded. According to (3), $\dot{\phi}=-nd^{n-1}\dot{d}-k\ddot{d}$. We ignored the subscript $i$ since it is assumed that there is only one obstacle. According to Assumption 1, we have the state space is bounded, thus both $d$ and $\dot{d}$ are bounded, which implies that $nd^{n-1}\dot{d}$ is bounded. Moreover, we have for all possible values of $a$ and $w$, there always exists a control $u$ to realize such $a$ and $w$ according to Assumption 1, which indicates the mapping from $u$ to $(a,w)$ is surjective. Since $a$ and $w$ are bounded and both can achieve zeros according to Assumption 1, we have $\forall u$, the corresponding $(a,w)$ are bounded. Since $\ddot{d}=-a\cos(\alpha)+v\sin(\alpha)w$, then $\ddot{d}$ is bounded. Hence $\mathcal{U}_{S}^{D}(x)=\mathcal{U}$ any $x$ that $\phi(x)-\eta<0$ and the claim is true. ∎
5.1.2 Proof of Proposition 1

### v2

@@ 4. Implicit Safe Set Algorithm for Systems with Negligible Sampling Time
4.1. Synthesize Safety Index for Continuous-time System
@@
These assumptions are easy to meet in practice. The first assumption ensures that the discrete-time approximation error is negligible, i.e., the system can essentially be treated as a continuous-time system. The second assumption makes the safety index design rule applicable with multiple moving obstacles.
Following the rules in (liu2014control, 25), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$,
(3) $\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i},$
where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$. We illustrate the behavior of the safety index $\phi$ with respect to $d$ and $\dot{d}$ under varying $k$ and $n$ in Figure 2. Our goal is to choose these parameters such that $\mathcal{U}_{S}^{D}(x)$ is always nonempty. Under the above assumptions, the safety index design rule is constructed below. The main idea of designing the safety index is to derive the sufficient condition of parameters of $\phi$ to ensure there exist safe controls for all states.
(a) k=1,n=2k=1,\ n=2
(b) k=1,n=6k=1,\ n=6
(c) k=10,n=2k=10,\ n=2
Figure 2. 3D heat map illustration of the behavior of the safety index $\phi$ with respect to $d$ and $\dot{d}$ under varying $k$ and $n$, where z axis denotes the value of $\phi$. The brighter color denotes higher $\phi$ value. In the figure axes, “+” denotes positive values and “–” denotes negative values.
@@
(4) $\frac{n(\sigma+d_{min}^{n}+kv_{max})^{\frac{n-1}{n}}}{k}\leq\frac{-a_{min}}{v_{max}},$
where $v_{max}$ is the maximum relative velocity that the robot can achieve in the obstacle frame. This design rule, together with the conditions in 1 and 2, guarantees that the safe control set $\mathcal{U}_{S}^{D}(x)$ is non-empty at any arbitrary state, which is the key to establishing Theorem 1. The proof of this non-emptiness is provided in Proposition 2.
4.2. Understanding of Safety Index Design Rule
Design philosophy. We treat $n$ as a user-chosen shaping parameter for the safety index, while $k$ is then selected to satisfy (4) so that a safe control set $\mathcal{U}_{S}^{D}(x)$ is non-empty for all states. We begin with the role of $k$.
Effect of $k$.
With safety index parameterization form in (3), the coefficient $k$ scales the index’s sensitivity to the relative speed $\dot{d}$. Increasing $k$ steepens the slope of $\phi$ along the $\dot{d}$ axis, so $\phi$ rises more aggressively when the robot is closing in (i.e., $\dot{d}<0$). Visually, compare Fig. 2(a) (k=1,n=2k{=}1,n{=}2) and Fig. 2(c) (k=10,n=2k{=}10,n{=}2): the surface tilts more strongly with $\dot{d}$, indicating earlier and stronger reactions to robot-obstacle approach speed ($\dot{d}$). Equation (4) explains why $k$ cannot be arbitrarily small: to keep the safe control set nonempty, $k$ must grow when the maximum relative velocity $v_{\max}$ is larger or when the available maximum deceleration $|a_{\min}|$ is smaller. Intuitively, faster environments or weaker braking demand earlier reactions, i.e., the larger the $k$ is, the more sensitive the safety mechanism is to $\dot{d}$, working like a time-to-collision style constraint.
Effect of $n$ (distance–velocity trade-off).
The distance term $d_{\min}^{n}-d^{n}$ in (3) sharpens as $n$ increases, redistributing attention toward short distances. Comparing Fig. 2(a) (k=1,n=2k{=}1,n{=}2) and Fig. 2(b) (k=1,n=6k{=}1,n{=}6), the surface rises much more quickly as $d$ shrinks. Thus, for the same $(d,\dot{d})$, a larger $n$ yields a larger $\phi$ when the robot is close, i.e., less tolerance to running toward an obstacle at short range, yet allows more tolerance to aggressive motion when the robot is far away. In other words, increasing $n$ changes the preference of the safety index: strict when close, permissive when far.
Why $n$ and $k$ should change together.
Once $n$ is chosen to set that preference, $k$ must be tuned to preserve feasibility under (4). A larger $n$ makes $\phi$ grow rapidly at small $d$; with bounded deceleration $|a_{\min}|$, there is only so much (and so fast) the controller can slow down. To ensure nonempty set of safe control that prevents $\phi$ from diverging upward, the design must respond more sensitively when $\dot{d}$ decreases, i.e., increase $k$. This coupling is explicit in (4), where $n$ and $k$ appear together (through a factor of the form $\tfrac{n}{k}\bigl{(}\sigma+d_{\min}^{n}+k\,v_{\max}\bigr{)}^{(n-1)/n}$): raising $n$ typically necessitates raising $k$ to keep the inequality satisfied. Practically: choose $n$ to sculpt proximity sensitivity; then increase $k$ as needed to guarantee a nonempty safe control set by considering speed/braking limits of the environment.
4.3. Sample-Efficient Black-Box Constrained Optimization
@@
(5) $\begin{split}&\min_{u_{t}\in\mathcal{U}}\|u_{t}-u^{r}_{t}\|^{2}\\ &\text{s.t. }\phi(f(x_{t},u_{t}))\leq\max\{\phi(x_{t})-\eta,0\}\end{split}$
A key insight we have is that: since the objective of (5) is convex, its optimal solution will always lie on the boundary of $\mathcal{U}_{S}^{D}(x)$ if $u^{r}\notin\mathcal{U}_{S}^{D}(x)$. Therefore, it is desired to have an efficient algorithm to find the safe controls on the boundary of $\mathcal{U}_{S}^{D}(x)$. To efficiently perform this black-box optimization, we propose a sample-efficient boundary approximation algorithm called Adaptive Momentum Boundary Approximation Algorithm (AdamBA), which is summarized in Algorithm 1. AdamBA leverages the idea of adaptive line search with exponential outreach/decay to efficiently search for the boundary points. The inputs for AdamBA are the approximation error bound ($\epsilon$), learning rate ($\beta$), reference control ($u^{r}$), gradient vector covariance ($\Sigma$), gradient vector number ($n$), reference gradient vector ($\vec{v}^{r}$), safety status of reference control ($S$), and the desired safety status of control solution ($S_{goal}$).
We illustrate the main boundary approximation procedures of AdamBA in Figure 3, where AdamBA is supposed to find the boundary points of $\mathcal{U}_{S}^{D}(x)$ (green area) with respect to the reference control $u^{r}\notin\mathcal{U}_{S}^{D}(x)$ (red star). The core idea of the AdamBA algorithm follows the adaptive line search (armijo1966minimization, 4), where three main procedures are included. I. AdamBA first initializes several unit gradient vectors (green vectors) to be the sampling directions, as shown in Figure 3(a). II. AdamBA enters the exponential outreach phase by exponentially increasing the gradient vector length until they reach $\mathcal{U}_{S}^{D}(x)$ as shown in Figure 3(b). Note that we discard those gradient vectors that go out of control space (red vectors). III. Next, AdamBA enters the exponential decay phase by iteratively applying Bisection search to find boundary points as shown in Figure 3(c). Finally, a set of boundary points will be returned after AdamBA converges as shown in Figure 3(d). Note that AdamBA and the line search methods are fundamentally similar to each other, while the purpose of AdamBA is to find the boundary of safe/unsafe action, while the line search methods are to find the minimum of a function. It can be shown that the boundary approximation error can be upper bounded within an arbitrary resolution.
(a) Initial gradient vectors
@@
Figure 3. Illustration of the procedure of the AdamBA algorithm. The algorithm is divided into four stages.
@@
Relation to momentum methods.
The proposed AdamBA algorithm employs a momentum-inspired strategy in step-size control. During the exponential outreach phase, the step length $\beta$ is geometrically amplified ($\beta\leftarrow 2\beta$) as long as the sampled point remains within the admissible region and retains its nominal safety label. This behavior emulates the inertial buildup of momentum and allows for rapid traversal of benign interior space. Upon detecting a change in the safety status, the algorithm transitions to a bisection-based exponential decay stage, refining the bracketing pair until a prescribed tolerance $\varepsilon$ is satisfied. This mirrors the damping phase observed in line-search corrections.
The interplay between these phases underscores the importance of hyperparameter selection. An initial $\beta$ that is too small results in wasted outreach doublings, while an overly aggressive choice may cause immediate overshoot and necessitate excessive corrective iterations. Similarly, the tolerance $\varepsilon$ governs the trade-off between computational cost and boundary fidelity. Empirically, setting $\beta$ to 5–10% of the state-space scale, using a growth factor of 2, and selecting $\varepsilon$ commensurate with actuator resolution have been found to minimize the total number of evaluations. Additionally, periodic resampling of probe directions helps mitigate directional bias in non-stationary environments. Collectively, these design principles enable AdamBA to harness momentum-like dynamics for fast yet precise boundary localization, making it a practical tool for safety-critical control applications.
4.4. Implicit Safe Set Algorithm for Continuous-Time System
@@
Leveraging AdamBA and the safety index design rule, we construct the implicit safe set algorithm (ISSA) to safeguard the potentially unsafe reference control. The proposed ISSA is summarized in Algorithm 2. Note that ISSA is presented under the context of discrete-time system with negligible sampling time, and $\mathcal{U}_{S}^{D}$ (line 3 of Algorithm 2) is the same as $\mathcal{U}_{S}$ when the sampling time is negligible. The key idea of ISSA is to use AdamBA for efficient search and grid sampling for worst cases. The inputs for ISSA are the approximation error bound ($\epsilon$), the learning rate ($\beta$), gradient vector covariance ($\Sigma$), gradient vector number ($n$) and reference unsafe control ($u^{r}$). Note that ISSA is only needed when the reference control is not safe.
ISSA contains an offline stage and an online stage. In the offline stage, we synthesize the safety index according to the design rules. There are two major phases in the online stage for solving (5). In online-phase 1, we directly use AdamBA to find the safe controls on the boundary of $\mathcal{U}_{S}^{D}(x)$, and choose the control with minimum deviation from the reference control as the final output. In the case that no safe control is returned in online-phase 1 due to sparse sampling, online-phase 2 is activated. We uniformly sample the control space and deploy AdamBA again on these samples to find the safe control on the boundary of $\mathcal{U}_{S}^{D}(x)$. It can be shown in Section 5.2 that the ISSA algorithm is guaranteed to find a feasible solution of (5) since the nonempty set of safe control is guaranteed by the safety index design rule in (4). And we show in Section 5 that the solution ensures forward invariance in the set $\mathcal{S}=\mathcal{X}_{S}\cap\mathcal{X}_{S}^{D}$. Note that Phase 2 may consume significant computational resources, potentially compromising real-time performance. To avoid jerky movements caused by delayed computations, parallel processing and other acceleration techniques can be employed to speed up grid sampling.
Although the ISSA algorithm builds upon the safe set algorithm (liu2014control, 25), the proposed safety index synthesis and AdamBA algorithm can be applied to other energy function-based methods to generate safe controls with or without an explicit analytical dynamics model.
5. Theoretical Results for ISSA When the Sampling Time is Negligible
@@
If the control system satisfies the conditions in 1 and 2, and the safety index design follows the rule described in Section 4.1, the implicit safe set algorithm in Algorithm 2 guarantees the forward invariance to the set $\mathcal{S}\subseteq\mathcal{X}_{S}$.
@@
5.1. Feasibility of the Safety Index for Continuous-Time System
@@
If 1) the dynamic system satisfies the conditions in 1 and 2; and 2) the safety index is designed according to the rule in Section 4.1, then the robot system in 2D plane has nonempty set of safe control at any state, i.e., $\mathcal{U}_{S}^{D}(x)\neq\emptyset,\forall x$.
@@ 5.1.1. Preliminary Results
If the dynamic system satisfies the conditions in 1 and 2 and there is only one obstacle in the environment, then the safety index design rule in Section 4.1 ensures that $\mathcal{U}_{S}^{D}(x)\neq\emptyset$ for $x$ such that $\phi(x)-\eta\geq 0$.
@@
(6) $\mathcal{U}_{S}^{D}(x)=\{u\in\mathcal{U}\mid\phi(f(x,u))\leq\phi(x)-\eta\}$
According to the first condition in 2, we have $dt\rightarrow 0$. Therefore, the discrete-time approximation error approaches zero, i.e., $\phi(f(x,u))=\phi(x)+dt\cdot\dot{\phi}(x,u)+\Delta$, where $\Delta\rightarrow 0$. Then we can rewrite (6) as:
(7) $\mathcal{U}_{S}^{D}(x)=\{u\in\mathcal{U}\mid\dot{\phi}\leq-\eta/dt\}$
@@
(8) $\forall x\text{ s.t. }\phi(x)\geq\eta,\exists u,\text{s.t. }\ddot{d}\geq\frac{\eta/dt-nd^{n-1}\dot{d}}{k}.$
Note that in the 2D problem, $\ddot{d}=-a\cos(\alpha)+v\sin{(\alpha)}w$ and $\dot{d}=-v\cos(\alpha)$. According to 1, there is a surjection from $u$ to $(a,w)\in W:=\{(a,w)\mid a_{min}\leq a\leq a_{max},w_{min}\leq w\leq w_{max}\}$. Moreover, according to (3), $\phi$ for the 2D problem only depends on $\alpha$, $v$, and $d$. Hence condition $\forall x\text{ s.t. }\phi(x)\geq\eta$ can be translated to $\forall(\alpha,v,d)\text{ s.t. }\sigma+d_{min}^{n}-d^{n}-kv\cos(\alpha)\geq\eta$. Denote the later set as
(9) $\Phi:=\{(\alpha,v,d)\mid\sigma+d_{min}^{n}-d^{n}-kv\cos(\alpha)\geq\eta,v\in[0,v_{max}],d\geq 0,\alpha\in[0,2\pi)\}.$
@@
(10) $\forall(\alpha,v,d)\in\Phi,\exists(a,w)\in W,\text{ s.t. }-a\cos(\alpha)+v\sin(\alpha)w$ $\geq\frac{\eta/dt+nd^{n-1}v\cos(\alpha)}{k}.$
@@
(11) $\forall(\alpha,v,d)\in\Phi,\exists(a,w)\in W,\text{ s.t. }-\frac{a}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k}\geq 0,$
@@
(12) $\min_{(\alpha,v,d)\in\Phi}\max_{(a,w)\in W}(-\frac{a}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k})\geq 0.$
@@
(13) $\min_{(\alpha,v,d)\in\Phi}\max_{(a,w)\in W}(-\frac{a}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k})$
(14) $=$ $\min_{(\alpha,v,d)\in\Phi}(-\frac{a_{min}}{v}+[\tan(\alpha)]_{+}w_{max}+[\tan(\alpha)]_{-}w_{min}-\frac{nd^{n-1}}{k})$
(15) $=$ $\min_{\alpha\in[0,2\pi),v\in(0,v_{max}]}(-\frac{a_{min}}{v}+[\tan(\alpha)]_{+}w_{max}+[\tan(\alpha)]_{-}w_{min}-\frac{n(\sigma+d_{min}^{n}+kv\cos(\alpha))^{\frac{n-1}{n}}}{k})$
(16) $=$ $-\frac{a_{min}}{v_{max}}-\frac{n(\sigma+d_{min}^{n}+kv_{max})^{\frac{n-1}{n}}}{k}.$
The first equality eliminates the inner maximization where $\tan(\alpha)]_{+}:=\max\{\tan(\alpha),0\}$ and $\tan(\alpha)]_{-}:=\min\{\tan(\alpha),0\}$. The second equality eliminates $d$ according to the constraint in $\Phi$. The third equality is achieved when $\alpha=0$ and $v=v_{max}$. According to the safety index design rule in Section 4.1, (16) is greater than or equal to zero. Hence (12) holds, which then implies that the inequality in (10) holds.
@@
If the dynamic system satisfies the contidions in 1 and 2 and there is only one obstacle in the environment, then the safety index design rule in Section 4.1 ensures that $\mathcal{U}_{S}^{D}(x)=\mathcal{U}$ for any $x$ that $\phi(x)-\eta<0$.
@@
(17) $\mathcal{U}_{S}^{D}(x)=\{u\in\mathcal{U}\mid\phi(f(x,u))\leq 0\}$
According to the first assumption in 2, we have $dt\rightarrow 0$. Therefore, the discrete-time approximation error approaches zero, i.e., $\phi(f(x,u))=\phi(x)+dt\cdot\dot{\phi}(x,u)+\Delta$, where $\Delta\rightarrow 0$. Then we can rewrite (17) as:
(18) $\mathcal{U}_{S}^{D}(x)=\{u\in\mathcal{U}\mid\dot{\phi}\leq-\phi/dt\}$
Note that $\eta=0$ according to the safety index design rule, then $\phi(x)-\eta<0$ implies that $\phi(x)<0$. Hence $\phi/dt\rightarrow\infty$ since $dt\rightarrow 0$. Then as long as $\dot{\phi}$ is bounded, $\mathcal{U}_{S}^{D}(x)=\mathcal{U}$.
Now we show that $\dot{\phi}$ is bounded. According to (3), $\dot{\phi}=-nd^{n-1}\dot{d}-k\ddot{d}$. We ignored the subscript $i$ since it is assumed that there is only one obstacle. According to 1, we have the state space is bounded, thus both $d$ and $\dot{d}$ are bounded, which implies that $nd^{n-1}\dot{d}$ is bounded. Moreover, we have for all possible values of $a$ and $w$, there always exists a control $u$ to realize such $a$ and $w$ according to 1, which indicates the mapping from $u$ to $(a,w)$ is surjective. Since $a$ and $w$ are bounded and both can achieve zeros according to 1, we have $\forall u$, the corresponding $(a,w)$ are bounded. Since $\ddot{d}=-a\cos(\alpha)+v\sin(\alpha)w$, then $\ddot{d}$ is bounded. Hence $\mathcal{U}_{S}^{D}(x)=\mathcal{U}$ any $x$ that $\phi(x)-\eta<0$ and the claim is true. ∎
5.1.2. Proof of Proposition 2